Calculate the mole fraction of HNO3 and water in solution of nitric acid in water containing 45percent of HNO3 by weight
1
Expert's answer
2020-04-13T02:35:03-0400

"\alpha(HNO_3)=[w(HNO_3)/M(HNO_3)] :[w(HNO_3)/M(HNO_3)+w(H_2O)/M(H_2O)]"

"\u03b1(HNO3\u200b)=[45\/63]:[45\/63+55\/18] =45\/63:(45\/63+55\/18)=0.199"

=19,9% mole fraction